A table top measures 2 m by 1 m 50 cm. What is its area in square metres?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
l = 2m
b = 1m 50 cm = 1.50 m
Area = l × b = 2 × 1.50
​= 3m^2

Thanks